OpenStudy (anonymous):

What is the length of st if S is at (-5,1) and T is at(-1,12) ? Round to the nearest hundredth. A,11.70.......B,12.53.......C,13.60.....D.14.32

OpenStudy (zzr0ck3r):

length of \[(x_0,y_0)and(x_1,y_1)\]is given by \[\sqrt{(x_1-x_0)^2+(y_1-y_0)^2}\]

OpenStudy (zzr0ck3r):

so\[\sqrt{(-1-(-5))^2+(12-1)^2}=\sqrt{4^2+11^2}=\sqrt{137}\approx11.70\]

OpenStudy (zzr0ck3r):

I hope you can do any ones like this in the future

OpenStudy (anonymous):

ok thx